OpenStudy (anonymous):

inverse of f(x)=x^3+9x+7

hero (hero):

There's two ways to do this

hero (hero):

Actually, satellite might want to help you. He's good at this

OpenStudy (anonymous):

i have no idea how to solve i cubic equation. do you?

hero (hero):

You just find the inverse...

OpenStudy (anonymous):

there is a formula but it is tremendously ugly. i was waiting to see the answer

hero (hero):

I've done it before, yes

OpenStudy (anonymous):

so idea is you want to solve \[y=x^3+9x+7\] for x or solve \[x=y^3+9y+7\] for y whatever you like. suppose y = 5 how would you do it?

hero (hero):

give me a minute or two

OpenStudy (anonymous):

are you sure the question isn't find \[f^{-1}(2)\] or something more manageable?

OpenStudy (anonymous):

F inverse of (7)

hero (hero):

hmmm

OpenStudy (anonymous):

aaah that is what i though! that we can do

OpenStudy (anonymous):

@hero take a look at this http://www.wolframalpha.com/input/?i=10%3Dx^3%2Bx%2B7 that was for 10. go to exact form and look at how ugly it is. i didn't think this was easy with a cube.

OpenStudy (anonymous):

but with a number we have hope. what this is really asking is "solve for x \[x^3+9x+7=7\] and that is easy enough

hero (hero):

I see

OpenStudy (anonymous):

right big ugly mess yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but with \[f^{-1}(7)\] we can do it as long as we understand that it just means find the value of x for which \[f(x)=7\]

OpenStudy (anonymous):

so put \[x^3+9x+7=7\] \[x^3+9x=0\] \[x(x^2+9)=0\] \[x=0\] done. and you probably could have done it with your eyeballs since the constant is 7

OpenStudy (anonymous):

so since \[f(0)=7\] we know \[f^{-1}(7)=0\]

hero (hero):

I was trying to find the inverse, but ran into a few problems

OpenStudy (anonymous):

i am sure you did! me too

hero (hero):

Satellite is it possible to go past level 100?

OpenStudy (anonymous):

thank you i was confused

hero (hero):

Here's your inverse, courtesy of wolfram alpha: http://www.wolframalpha.com/input/?i=solve%28x%3Dy^3%2By%2B7%2C+y%29